Hebe attempts a crossword puzzle every day. The number of puzzles she completes in a week (7 days) is denoted by X.
- (i)State two conditions that are required for X to have a binomial distribution.[2]
- (ii)On average, Hebe completes 7 out of 10 of these puzzles. Use a binomial distribution to find the probability that Hebe completes at least 5 puzzles in a week.[3]
- (iii)Use a binomial distribution to find the probability that, over the next 10 weeks, Hebe completes 4 or fewer puzzles in exactly 3 of the 10 weeks.[3]
Mark scheme
Marking points from the official mark scheme (May/June 2017). The mark or mark type is shown in brackets.
- [B1] constant probability (of completing) Any one condition of these two
- [B1] independent trials/events The other condition
- [M1] P(5, 6, 7) = 7C5(0.7)5(0.3)² + 7C6(0.7)6(0.3)¹ + (0.7)7 Bin term 7Cₓ(0.7)ˣ(0.3)⁷⁻ˣ, x ≠ 0, 7
- [A1] Correct unsimplified answer (sum) OE
- [A1] = 0.647
- [M1] P(0, 1, 2, 3, 4) = 1 – their ‘0.6471' = 0.3529 Find P( ≤4) either by subtracting their (ii) from 1 or from adding Probs of 0,1,2,3,4 with n=7 (or 10) and p = 0.7
- [M1] P(3) = 10C3(0.3529)3(0.6471)7 10C3 (their 0.353)³(1 – their 0.353)⁷ on its own
- [A1] = 0.251
Write your answer and get it marked
On Oracle Prep you can write or photograph your own answer and have it marked by AI against this mark scheme, with feedback on what you missed.
Sign up free →